One number is 3 more than half of a second number. If the average of the two numbers is 15 1/2, find the smaller of the two numbers.
Found 2 solutions by checkley71, wgomero:
Answer by checkley71(8403)

(X+Y)/2=15.5 THEN X+Y=31 OR X=31-Y AND X=3+Y/2 THUS 31-Y=3+Y/2 OR Y+Y/2=31-3 OR
3/2Y=28 OR 3Y=28*2 OR 3Y=56 OR Y=18 2/3 THEN X=31-18/2/3 OR X=12 1/3 [THE SMALLER NUMBER]
PROOF 18 2/3+12 1/3=31 OR 31=31 AND 12 1/3=3+(18 2/3)/2 OR 12 1/3=3+9 1/3 OR
12 1/3=12 1/3

Answer by wgomero(28)

resolution:
let the second number be x
the other number will be 3 + x/2
average of the two numbers is (x + (3 + x/2) )/ 2= 15 1/2;
(x + 3 + x/2)/2 = 31/2; since the equation has the same denominator cross multiple by 2 then remove 2 from both sides and solve the linear equestion that contains fraction; the linear equation will be: 2x + 6 + x = 62
solve by x: 3x = 56; x = 18.66666; 18.7
the other number will be 3 + (18.7)/2 = 12.35
add numbers (18.7 + 12.35) = 31.05; average 31.05 /2 = 15.5
the small number is 12.35